Calculate the pressure exerted on an elevator floor which is 2 meters length and 3 meter wide, if 2 people whose combined mass is 160 kg are standing on it. (g = 10 m/s2)
1
Expert's answer
2020-12-23T10:19:44-0500

In general, if the force is evenly distributed over the entire area, then



"p=\frac{F}{A}=\frac{160\cdot10}{2\cdot3}=267(Pa)"